The following table gives data on the Boston Red Sox wins and runs. Construct a scatterplot...

  1. The following table gives data on the Boston Red Sox wins and runs. Construct a scatterplot with runs on the horizontal axis and wins on the vertical axis.

a) Do you think there is a linear relationship between the number of runs and the number of wins in a given year?

b) The linear regression line for the data is: Wins = 24.5+ 0.08 Runs. Use this model to predict the number of wins the Red Socks would have in a season where they had 800 Runs. (Round your answer to one decimal place)

c) What is the slope of the regression line? Interpret the slope.

d) What is the y-intercept of the regression line? Can you interpret the y-intercept? Why or why not? If you can, do so.

RUNS

WINS

872

95

845

95

867

96

820

86

910

95

949

98

961

95

859

93

772

82

792

85

836

94

876

92

851

78

928

85

791

86

552

54

686

80

599

73

731

84

699

88
Answer #1

The given information is related to Boston Red Sox wins and runs. Follow the below steps to construct the scatter plot usingWins 0 200 400 600 800 1000 1200 Runs(a) Yes, there is a linear relationship exists between the number of runs and the number of wins in a given year. From the ScWins = 24.5+0.08 Runs Wins = 24.5+0.08(800) Wins = 88.5 Therefore, the number of wins the Red Socks would have in a season whThe slope of the regression line is 0.08. Interpretation: If one unit change in Runs 0.08 times change will be occurred in Wi